Qn 5. In a scalene triangle whose sides have lengths a, b and c, consider

the bisector r of the angle formed by a and b. Compute, providing your

working based on Euclidean geometry, the ratio between the lengths of the

two segments that the bisector r determines on the side of length c when

intersecting it.


1
Expert's answer
2022-03-01T04:21:51-0500

Let ABC be the triangle. Here AB = c, AC = b, BC = a. We look at the bisector of the angle C, let us call this bisector CM = r. We should determine the ratio AM:MB.

Let us denote the half of angle C as "\alpha." So the area of triangle ACM is "\frac12 AC\cdot CM\cdot \sin\alpha" and area of triangle BCM is "\frac12 BC\cdot CM\cdot \sin\alpha" . Therefore, the ratio of areas is AC:BC. But if we write the formula for area using the height and side, we'll obtain the area of ACM to be "\frac12\cdot AM\cdot h," area of BCM to be "\frac12\cdot BM\cdot h," so the ratio of areas is AM:BM.


Finally, we obtain AC:BC = AM:BM, or b:a = AM:BM. We also know that AM+BM = c, so "AM = c\cdot\frac{b}{a+b}, BM = c\cdot\frac{a}{a+b}" .